The enthalpy of formation of NH3(g) as per the following reaction is -46.11 kJ/mol 

at 298 K. NE(g) + 3H2(g) ----> 2NH3(g)

 Calculate the value of enthalpy of formation of NH2(g) at 100°C. The Cp(J/K/mol)

 

values are given as: N2(g) = 29.12; H2(g) = 28.82; NH3(g) = 35.06. Assume the 

given heat capacity values to be temperature independent in the range. 


1
Expert's answer
2021-04-01T06:01:48-0400

Solution:

The balanced chemical equation:

N2(g) + 3H2(g) → 2NH3(g), ΔrHo = -46.11 kJ/mol

Assume heat capacities are independent of temp T.

Thus,

ΔrCp = ∑Cp(products) − ∑Cp(reactants)

ΔrCp = 2×Cp(NH3) - (3×Cp(H2) + Cp(N2))

ΔrCp = (2 × 35.06) - (3 × 28.82 + 29.12) = -45.46 J/K/mol


The temperature dependence of enthalpy: dH = CpdT (at constant pressure).

And so for a temperature change from T1 to T2:



If Cp is independent of temperature, then:

ΔrH(T2) = ΔrH(T1) + ΔrCp × (T2 - T1)


ΔrCp = -45.46 J/K/mol

T1 = 298 K

T2 = 100°C + 273 = 373 K

Hence,

ΔrH at 100°C = -46110 J/mol + (-45.46 J/K/mol) × (373 K - 298 K) = -49519.5 J/mol = -49.52 kJ/mol

ΔrH at 100°C = -49.52 kJ/mol


Answer: ΔrH at 100°C = -49.52 kJ/mol.